OpenStudy (anonymous):

Solve the equation for y. 8x – 9y = 11

OpenStudy (anonymous):

is it multiple choice?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

what are the answer choices?

OpenStudy (anonymous):

A. y= -9/8x-9/11 B. y=9/8x + 11/8 C. y=-8x -11 D. y= y = 8/9x - 11/9

OpenStudy (anonymous):

ok give me a couple min

OpenStudy (anonymous):

tyt

OpenStudy (anonymous):

First you move the 8x because you have to leave the 9y alone 8x-9y=11 -8x -8x -9y=11-8x Now you have to move the -9, so the inverse of multiplying is dividing so you divide -9. -9y/-9= 11/-9 -8x/-9 since you can't have negative as a denominator it would be: y= -11/9+8x/9 You have to change the signs

OpenStudy (anonymous):

does it make sense?